40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 10 14:58:01 2009, Jenne, Update, PEM, 2nd set of Guralp channels plugged into ADCU 
    Reply  Mon Aug 10 18:12:25 2009, Jenne, Update, PEM, 2nd set of Guralp channels plugged into ADCU 
Message ID: 1882     Entry time: Mon Aug 10 18:12:25 2009     In reply to: 1872
Author: Jenne 
Type: Update 
Category: PEM 
Subject: 2nd set of Guralp channels plugged into ADCU 

Quote:

The second set of Guralp channels is now plugged into the PEM ADCU, into channels which are confirmed to be working.  (Method: 1Vpp sine wave into channel, check with DataViewer).

 

Direction, Channel Name, .ini chnum, BNC plug # on ADCU

Vertical: C1:PEM-SEIS_GUR_VERT, 15023, #24

N/S (should be Y when the seismometer is put in place): C1:PEM-TEMP_2, 15001, #2

E/W (should be X when the seismometer is put in place): C1:PEM-TEMP_3, 15002, #3

 

There is IFO work going on, so I don't want to rename the channels / restart fb40m until a little later, so I'll just use the old TEMP channel names for now. 

 

There is something totally wrong with the E/W channel.  I can look at all 3 channels on a 'scope (while it's on battery, so the op-amps in the breakout box aren't grounded), and VERT and NS look fine, and when I jump around ("seismic testing"), they show spikes.  But the EW channel's signal on the 'scope is way smaller, and it doesn't show anything when I jump. 

 

I might use the handheld Guralp tester breakout box to check the seismometer.  Also, a suspicion I have is that whoever put the box back in on Friday night after our final noise measurements left the inputs shorted for this one channel.  It's the 3rd channel in the set, so it would be most likely to be stuck shorted...  Investigations will ensue.

 All the channels are now good, and all the names are back to making sense. 

The problem with EW2 was in fact that the alligator clip used to short the inputs during the noise test Friday night was left in the box.  Not great, but now it's taken care of, and we have recorded data of the noise of the breakout box, so we can include that in our plots to see if we're at the limit of how good we can do at subtracting noise.

 

The channels are now named thusly:

C1:PEM-SEIS_GUR_VERT  (BNC input #24, .ini channel #15023)

C1:PEM-SEIS_GUR_EW     (BNC input #3, .ini channel #15002)

C1:PEM-SEIS_GUR_NS      (BNC input #2, .ini channel #15001)

C1:PEM-SEIS_MC1_X         (BNC input #11, .ini channel #15010)

C1:PEM-SEIS_MC1_Y        (BNC input #12, .ini channel #15011)

C1:PEM-SEIS_MC1_Z       (BNC input #10, .ini channel #15009)

C1:PEM-SEIS_MC2_Y (Ranger, which for the Huddle Test is oriented VERTICALLY)   (BNC input #4, .ini channel #15003)

 

Now we wait.....and tomorrow extract the noise of each of the seismometers from this!

 

 

ELOG V3.1.3-